A small electric car has a max constant acceleration of 1m/s^2 a max constant deacceleration of 2m/sec2
and a max speed of 20m/sec.What is the total  amount of time it would take to drive this car for 1km starting from rest ?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Time it takes to reach 20m/s.
v=u+at
=>t=20s

S=at2/2
=>S=400m

Time it takes to reach 0 again because it will start to decelerate
v=u+at
=>20=2t
=>t=10

Distance it will travel in this 10s=ut+at2/2 = 20x10-2x100/2
=100m

Total distance=500m time taken for it 30s
for distance 1000m time will be 60s since the same process will repeat again.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)